OpenStudy (anonymous):

simplify sqrt 300x^16

OpenStudy (pfenn1):

\[\sqrt{300x^{16}}=\sqrt{2 \times 2 \times 3 \times 5 \times 5 \times x^8 \times x^8}\]Does that give you any hints?

OpenStudy (anonymous):

yeah, a little. I think I can figure it out from here. Thanks

OpenStudy (pfenn1):

you're welcome